• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name suleiman

    Log entry time 08:04:21 on December 5,2006

    Entry number 190093

    keyword=Owl Shift Summary



    00:06 Run 1318 He3.
    00:50 Reduced S1 Right Arm voltages to 60% of their original
    values.
    01:04 Run 1319 He3.

    01:35 Reduced S1 Right Arm voltages to 67% of their original
    values.
    01:42 Run 1320 He3.
    02:15 Run 1320 He3. New PS=1,6000,10,150,1
    02:50 Run 1320 He3. New PS=1,6000,10,150,1
    03:45 Run 1320 He3. New PS=1,6000,1,150,1

    04:00 Reduced S1 Right Arm voltages to 75% of their original
    values.
    04:07 Run 1324 He3. New PS=100,2000,10,150,1

    The protons are passing the S1-R ADC threshold and the accidental
    rate is big. Changed HV back to 0.67% of its original value
    04:43 Run 1325 He3. New PS=1,2000,10,150,1

    The threshold seems on the edge of the deuteron signal, go to
    70% HV.

    05:45 Reduced S1 Right Arm voltages to 70% of their original
    values.
    05:52 Run 1326 He3. New PS=10,2000,10,100,1
    06:52 Run 1327 He3. New PS=10,2000,10,100,1

    07:50 Move target to 20 cm Dummy
    08:00 Run 1328 He3. New PS=1,2000,1,100,1